OpenStudy (anonymous):

suppose in the early 1900s the levels of CO2 in the atmosphere was only 64% of the levels of CO2 today. what percentage more is the level of CO2 today compared to the levels of CO2 in the early 1900s?

OpenStudy (anonymous):

@perl

OpenStudy (perl):

there is a typo in your questions,

OpenStudy (perl):

i have a hard time answering when i see typos in questions

OpenStudy (perl):

suppose in the early 1900s the levels of CO2 in the atmosphere was on only 64% of the levels of CO2 today. was on?

OpenStudy (anonymous):

i fixed it /.\

OpenStudy (perl):

i think its just 100 - 64 % or 36% more

OpenStudy (perl):

i think its just (100 - 64) % or 36% more

OpenStudy (perl):

the question is a bit ambiguous, q

OpenStudy (tkhunny):

Yes, ambiguous. Another take. 1900 64 Today 100 Percent Increase: (100-64)/64 = 0.5625

OpenStudy (anonymous):

can you explain to me why you did that?

OpenStudy (tkhunny):

1) Where did it start? 2) How much did it change? 3) Calculate percent increase.

OpenStudy (perl):

I did that as my second approach

OpenStudy (perl):

the directions say that the CO2 level in 1900 was 64% of today. So we can assume, without loss of generality, that there are 100 pounds of CO2 in the atmosphere today. That means in 1900 there were 64 pounds of CO2 in the atmosphere.

OpenStudy (perl):

Now , what is the increase from 64 pounds to 100 pounds, in terms of percents?

OpenStudy (perl):

you can solve the equation 64 ( 1 + r ) = 100 , where r is the rate of increase in decimal

OpenStudy (perl):

r = 100 /64 - 1 . this will give you a decimal, which you can convert to a percent

OpenStudy (perl):

r = .5625 , which is 56.25% increase.